a desert outpost has a water supply that is sufficient to last 21 days for 15 people. at the same average rate of water consumption per person , how many days would the water supply last for 9 people?
Answer by stanbon(75887)

a desert outpost has a water supply that is sufficient to last 21 days for 15 people. at the same average rate of water consumption per person , how many days would the water supply last for 9 people?
----
# of days and # of people are inversely related:
d = k/n
---
Solve for "k" using "outpost has a water supply that is sufficient to last 21 days for 15 people"
21 = k/15
k = 15*21
-----
Equation:
d = 315/n
-------------
Solve "how many days would the water supply last for 9 people?"
d = 315/9
# of days = 35
=====================
Cheers,
Stan H.
===========